# The Harrow-Hassidim-Lloyd algorithm with qutrits

### v1

@@ V.2 Quantum chemistry simulations
We now focus on the equilibrium geometry obtained from our PEC (1.4 Bohr), to study the effect of $n_{r}$ on the correlation energy using qubit and qutrit HHL implementations. The results are provided in Figure 6, with Table A.3 of the Appendix presenting the accompanying data. We stop at $n_{r}=6$ in the figure due to computational limitations from qutrit HHL. We note that for the purposes of our statevector calculations, setting $n_{r}=6$ would result in $2^{6}=64$ multi-controlled rotation gates in the controlled-rotation module in the case of qubits, whereas for qutrits, we end up with $3^{6}=729$ multi-controlled rotation gates. This observation underscores the need to develop more computationally efficient classical computing routines for qudit HHL implementations in general, and we defer the exercise for a future study. We now turn our attention back to the results in Figure 6, which show that for a fixed number of qudits in the clock register, qutrit HHL performs much better than its qubit counterpart. This is not surprising, given that the eigenvalues are captured to $3^{n_{r}}$ precision for qutrits, whereas it is $2^{n_{r}}$ for qubits. We also see that ‘convergence’ to the expected correlation energy value is achieved faster for qutrit HHL than qubit HHL for the same reason. We recall that the HHL algorithm involves a constant $C$ (introduced in Section III.3). We have modified this constant as by taking its ternary expansion (for qutrits) and its binary expansion (for qubits) up to the same precision as is achieved by the number of qudits in the clock register. Figure 6 shows that the correlation energies have improved with the ternary and binary expanded $C$ values for qubit and qutrit HHL respectively.
@@
VI Comparison of qubit and qutrit HHL
VI.1 Analyzing the number of qudits
@@
    Clock register:
    Let the required precision in decimal numbers be $p$ digits. This means that we want the error to be $\epsilon=10^{-p}$. In the previous sections, the number of qudits in this register are denoted by $n_{r}$. Since the captured eigenvalues $|\phi\rangle$ are stored in these $n_{r}$ qudits as $|\tilde{\phi}\rangle$, $|\phi-\tilde{\phi}|=\frac{1}{d^{n_{r}}}$, where $d$ is the dimension of the qudit. In such a case, the maximum error $\epsilon$ will be $\frac{1}{d^{n_{r}}}$, and hence
@@
    State register:
    In order to encode the $N\crossproduct 1$ dimensional $\vec{b}$ in the state register, $m$ qudits are required such that $N\leq d^{m}$. Since we work in the binary (for qubits) and the ternary (for qutrits) frameworks, the size of $|b\rangle$ is accordingly $2^{m^{(b)}}$ or $3^{m^{(t)}}$. Here, $m^{(b)}$ and $m^{(t)}$ represent the number of qubits and qutrits in the state register respectively. The $\vec{b}$ must be encoded in these quantum states, and hence $m^{(b)}$} or $m^{(t)}$} are chosen such that they accommodate the $N\crossproduct 1$ dimensional $\vec{b}$, while the remaining $d^{m}-N$ entries are taken as zeroes. From this condition, it can be shown that
@@
    HHL ancilla:
    A single qudit is required for the HHL ancilla regardless of it being a qubit or a qutrit.
@@
VI.2 Analyzing the number of 2-qudit gates
The overall resource cost can be estimated by the number of two-qudit gates, since they are considered to be the major sources of error at the physical level. We do the counting module-wise as follows:
@@ VI.2 Analyzing the number of 2-qudit gates
    Controlled-unitaries within QPE:
    The total number of controlled-unitaries for qubits is $2^{{n_{r}}^{(b)}}-1$, while for qutrits it is $\frac{1}{2}(3^{{n_{r}}^{(t)}}-1)$. However, for a given precision $p$, using equation 5, we can show that
@@
    Therefore, the number of controlled-unitaries is halved for the qutrit case. We now assume a recursive decomposition scheme for the $m^{(b)}$ qubits in the state register as proposed in Reference [shende_decompo], and a similar decomposition scheme for the $m^{(t)}$ qutrits [jiang_qutrit_decompo], and work out both the cases to find that the number of two-qudit gates are of the same order. This is achieved because of the fact that for a given precision $m^{(b)}$ and $m^{(t)}$ are related using equation 7. We note that while the decomposition for qubits goes as $4^{m^{(b)}}$ and the one for qutrits scales as $9^{m^{(t)}}$, the number of two-qudit gates in a controlled-unitary gate scales quadratically in the $A$ matrix size due to amplitude encoding in both the cases.
@@
    IQFT module within QPE:
    The count for two-qudit gates in this module is $\frac{{{n_{r}}^{(b)}}^{2}-1}{2}$ for qubits and $\frac{{{n_{r}}^{(t)}}^{2}-1}{2}$ for qutrits, which can be simplified using equation 6 as $\approx\frac{0.39\;{{n_{r}}^{(b)}}^{2}-1}{2}$. Therefore, the IQFT module with qutrits has $\approx 39\%$ gates as its qubit counterpart.
@@
    Controlled-rotation module:
    The number of multi-controlled rotation gates for the controlled rotation module is $d^{n_{r}}$. This means that qutrit HHL requires $3^{{n_{r}}^{(t)}}$ multi-controlled rotations and qubit HHL requires $2^{{n_{r}}^{(b)}}$ multi-controlled rotations. Using equation 8, we can show that the number of multi-controlled rotation gates in the controlled-rotation module is equal for both the qubit and qutrit HHL implementations. The decomposition to two-qudit gates again goes as in the first point. Hence, the number of two-qudit gates in the controlled-rotation module are comparable for both the implementations.
VII Conclusion
In this work, we extend the HHL algorithm from its qubit form to the qutrit framework, by designing the circuit (including a simplified controlled-rotation module) as well as developing qubit and qutrit HHL codes in the BQSKit framework. To demonstrate the working of the qutrit HHL algorithm, we carry out toy matrix calculations as well as quantum chemical calculations to generate the potential energy curve of the model $H_{2}$ molecule. We obtain correlation energies to within 0.02 percent precision for the latter, and also benchmark our results with CISD and LCCSD calculations using classical methods.

### v2

@@ VI.2 Quantum chemistry simulations
We now focus on the equilibrium geometry obtained from our PEC (1.4 Bohr), to study the effect of $n_{r}$ on the correlation energy using qubit and qutrit HHL implementations. The results are provided in Figure 8, with Table A.3 of the Appendix presenting the accompanying data. We stop at $n_{r}=6$ in the figure due to computational limitations from qutrit HHL. We note that for the purposes of our statevector calculations, setting $n_{r}=6$ would result in $2^{6}=64$ multi-controlled rotation gates in the controlled-rotation module in the case of qubits, whereas for qutrits, we end up with $3^{6}=729$ multi-controlled rotation gates. This observation underscores the need to develop more computationally efficient classical computing routines for qudit HHL implementations in general, and we defer the exercise for a future study. We now turn our attention back to the results in Figure 8, which show that for a fixed number of qudits in the clock register, qutrit HHL performs much better than its qubit counterpart. This is not surprising, given that the eigenvalues are captured to $3^{n_{r}}$ precision for qutrits, whereas it is $2^{n_{r}}$ for qubits. We also see that ‘convergence’ to the expected correlation energy value is achieved faster for qutrit HHL than qubit HHL for the same reason. We recall that the HHL algorithm involves a constant $C$ (introduced in Section III.3). We have modified this constant by taking its ternary expansion (for qutrits) and its binary expansion (for qubits) up to the same precision as is achieved by the number of qudits in the clock register. Figure 8 shows that the correlation energies have improved with the ternary and binary expanded $C$ values for qubit and qutrit HHL respectively.
@@
Case 2: 2-qutrit input state
We also perform a simple statevector calculation for the case of two qutrits in the state register following our approach described in Section IV. Here, we consider the same problem of the $H_{2}$ molecule in the 631G6-31G basis, but this time without freezing any of the orbitals, thereby working in an active space of one occupied and three virtual molecular orbitals, which makes it a total of eight spin orbitals. Using the GUGA-CISD approach, we obtain a Hamiltonian matrix of dimension $6\times 6$, which on slicing according to the LCCSD method reduces to a $5\times 5$ AA matrix, which in turn is padded to $9\times 9$ and used as the input in our qutrit HHL implementation. For the purposes of this work, we do a proof-of-principle demonstration of the decomposition schemes only for the equilibrium geometry of $H_{2}$ and obtain results within $0.01\%$ precision as compared to the expected correlation energy from classical LCCSD calculations. The correlation energy obtained from qutrit HHL calculation with 5 clock registers is 0.025401-0.025401 Hartree, while the LCCSD reference value is 0.025527-0.025527 Hartree. Next, we move to an analysis of qubit and qutrit HHL implementations in terms of resources.
VII Comparison of qubit and qutrit HHLs
VII.1 Analyzing the number of qudits
@@
    Clock register:Let the required precision in decimal numbers be $p$ digits. This means that we want the error to be $\epsilon=10^{-p}$. In the previous sections, the number of qudits in this register are denoted by $n_{r}$. Since the captured eigenvalues $|\phi\rangle$ are stored in these $n_{r}$ qudits as $|\tilde{\phi}\rangle$, $|\phi-\tilde{\phi}|=\frac{1}{d^{n_{r}}}$, where $d$ is the dimension of the qudit. In such a case, the maximum error $\epsilon$ will be $\frac{1}{d^{n_{r}}}$, and hence
@@
    State register: In order to encode the $N\crossproduct 1$ dimensional $\vec{b}$ in the state register, $m$ qudits are required such that $N\leq d^{m}$. Since we work in the binary (for qubits) and the ternary (for qutrits) frameworks, the size of $|b\rangle$ is accordingly $2^{m^{(b)}}$ or $3^{m^{(t)}}$. Here, $m^{(b)}$ and $m^{(t)}$ represent the number of qubits and qutrits in the state register respectively. The $\vec{b}$ must be encoded in these quantum states, and hence $m^{(b)}$} or $m^{(t)}$} are chosen such that they accommodate the $N\crossproduct 1$ dimensional $\vec{b}$, while the remaining $d^{m}-N$ entries are taken as zeroes. From this condition, it can be shown that
@@
    HHL ancilla: A single qudit is required for the HHL ancilla regardless of it being a qubit or a qutrit.
@@
VII.2 Analyzing the number of 2-qudit gates
The overall resource cost can be estimated by the number of 2-qudit gates, since they are considered to be the major sources of error at the physical level. We do the counting module-wise as follows:
@@ VII.2 Analyzing the number of 2-qudit gates
    Controlled-unitaries within QPE: The total number of controlled-unitaries for qubits is $2^{{n_{r}}^{(b)}}-1$, while for qutrits it is $\frac{1}{2}(3^{{n_{r}}^{(t)}}-1)$. However, for a given precision $p$, using Equation (37), we can show that
@@
    Therefore, the number of controlled-unitaries is halved for the qutrit case. We now assume a recursive decomposition scheme for the $m^{(b)}$ qubits in the state register as proposed in Reference [shende_decompo], and a similar decomposition scheme for the $m^{(t)}$ qutrits [jiang_qutrit_decompo], and work out both the cases to find that the number of 2-qudit gates are of the same order. This is achieved because of the fact that for a given precision $m^{(b)}$ and $m^{(t)}$ are related using Equation (39). We note that while the decomposition for qubits goes as $4^{m^{(b)}}$ and the one for qutrits scales as $9^{m^{(t)}}$, the number of 2-qudit gates in a controlled-unitary gate scales quadratically in the $A$ matrix size due to amplitude encoding in both the cases.
@@
    IQFT module within QPE:The count for 2-qudit gates in this module is $\frac{{{n_{r}}^{(b)}}^{2}-1}{2}$ for qubits and $\frac{{{n_{r}}^{(t)}}^{2}-1}{2}$ for qutrits, which can be simplified using Equation (38) as $\approx\frac{0.39\;{{n_{r}}^{(b)}}^{2}-1}{2}$. Therefore, the IQFT module with qutrits has $\approx 39\%$ gates as its qubit counterpart.
@@
    Controlled-rotation module: The number of multi-controlled rotation gates for the controlled rotation module is $d^{n_{r}}$. This means that qutrit HHL requires $3^{{n_{r}}^{(t)}}$ multi-controlled rotations and qubit HHL requires $2^{{n_{r}}^{(b)}}$ multi-controlled rotations. Using Equation (40), we can show that the number of multi-controlled-rotation gates in the controlled-rotation module is equal for both the qubit and qutrit HHL implementations. The decomposition to 2-qudit gates again goes as in the first point. Hence, the number of 2-qudit gates in the controlled-rotation module are comparable for both the implementations.
We note at this point that counting the number of non-Cliffords for a fault-tolerant implementation is outside the scope of the current study.
VII.3 Gate counts for physical Hamiltonians and with Trotterization and Weyl-Heisenberg gadgets
[…]
In the previous section, we discussed the number of controlled-unitaries being $2^{{n_{r}}^{(b)}}-1$ and $\frac{1}{2}(2^{{n_{r}}^{(b)}}-1)$ (Equation (40)) for qubit and qutrit HHL implementations respectively. We now estimate the gate count for each such controlled-unitary. As discussed in Section IV, each term in the expansion of $e^{iAt}$ upon Trotterization (first order and first step) translates to a Pauli gadget (for qubits) and WH gadget (for qutrits). For qubits, the decomposition for the controlled-unitary yields $2$ Pauli gadgets each containing $2(m^{(b)}+1)$ 2-qubit gates. Using Equation (32), each controlled WH gadget is decomposed to $3$ WH gadgets each having $2(m^{(t)}+1)$ 2-qutrit gates. Comparing the cases, and considering the relation between $m^{(b)}$ and $m^{(t)}$ from Equation (39), we infer that the number of 2-qubit gates goes as $4(m^{(b)}+1).(poly(m^{(b)}))$, whereas the number of 2-qutrit gate count is $\frac{6}{2}(m^{(t)}+1).(Poly(m^{(t)}))$, which can be expressed in terms of m(b)m{(b)} as $3(0.63m^{(b)}+1).(Poly(m^{(t)}))$. We use ‘Poly’ for the qutrit case as opposed to ‘poly’ for the qubit case to distinguish between the two polynomial functions. Therefore, the number of gates between the qubit and qutrit cases are still comparable upon applying HHL to physical Hamiltonians and using Trotterization and relevant gadgets for decomposing matrix exponentials.
VIII Conclusion
In this work, we extend the HHL algorithm from its qubit form to the qutrit framework, by designing the circuit (including a simplified controlled-rotation module and finding Weyl-Heisenberg gadgets to implement the $e^{iAt}$ module of qutrit HHL), as well as developing qubit and qutrit HHL codes in the open-source BQSKit framework. To demonstrate the working of the qutrit HHL algorithm, we carry out toy matrix calculations as well as quantum chemical calculations involving 1-qutrit state registers to generate the potential energy curve of the model $H_{2}$ molecule. Furthermore, we make use of our gadgets to find the ground state energy in the equilibrium geometry of the same system, but with a 2-qutrit state register. We obtain correlation energies to within 0.02 percent precision for the 1-qutrit state example and to within 0.01 percent for the 2-qutrit state example (the equilibrium geometry case), and also benchmark our results with CISD and LCCSD calculations using classical methods.